# Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]

### v1

@@ 2.3. Motivation
The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field. A significant amount of works on GNN training do not compare or evaluate their systems against systems using a different training approach(full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021a; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021). In the following, we review 33 papers on GNN training that do make claims comparing systems in the two classes. We summarize the claims in these papers in Table 1.
@@
We classify the papers into three broad groups. The first group consists of papers that propose full-graph training. Out of the 17 papers in this group, 10 provide experimental evidence. The rest 7 papers make claims that full-graph training is superior in terms of higher accuracy, lower epoch time, and convergence guarantees and support the claims by only citing other papers. The second group includes 12 papers that propose mini-batch training. Among these papers, only 4 provide experimental evidence to support claims that full-graph training is difficult to scale and requires more memory, while mini-batch training has lower time-to-accuracy. The last group consists of 4 papers that do not propose one specific training approach. Overall, we find the following shortcomings in the existing literature, which motivate our comprehensive experimental comparison.
@@
Conflicting claims about how the two approaches compare. Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems. Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level. Other works claim full-graph training is impractical, requires high memory and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021). Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
@@
Comparing performance only in terms of epoch time. Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023). As we will show, a comparison in terms of time to accuracy is more informative given the differences in the two training approaches. Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
@@
Experiment design. Some works (Jia et al., 2020; Liu et al., 2021; Li et al., 2024; Kaler et al., 2022) do not always use the same GNN model (e.g., GraphSAGE or GAT) for the two approaches as a basis of comparison. This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique. Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023a, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022). As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa. Using hyperparameter tuning is necessary to perform a fair comparison.
@@
Table 2. Dataset statistics. (*) The input features and the training data for Orkut are synthetic
Datasets #Nodes #Edges #Classes #Features Train/Val/Test
Pubmed 19,717 108,365 3 500 18,217/500/1,000
Ogbn-arxiv 169,343 1,166,243 40 128 90,941/29,799/48,603
Reddit 232, 965 11, 606, 919 41 602 152,410/23,699/55,334
Ogbn-products 2,449,029 61,859,140 47 100 195,922/48,980/2,204,127
Orkut 3,072,441 117,185,083 2* 128* 1,843,464/614,488/614489*
Ogbn-papers100M 111,059,956 1,615,685,872 172 128 1,207,179/125,265/214,338
Table 3. Hyperparameter search space
Parameter range/values
Number of Hidden Layers [2, 10]
Size of hidden layers [16, 1024]
Aggregator (GraphSage only) { mean, gcn, pool }
Number of heads (GAT only) [1, 12]
Sampling Fanout (MB only) [4, 50]
Learning rate [0.0001, 0.01]
Dropout [0.2, 0.9]
Mini-batch size (MB only) [128, 4096]
3. Experimental setup
@@
Datasets. Our experiments use six datasets: pubmed (Namata et al., 2012), ogbn-arxiv (Hu et al., 2020), reddit (Hamilton et al., 2017), ogbn-products (Hu et al., 2020), orkut (Yang and Leskovec, 2012) and ogbn-papers100m (Hu et al., 2020) (see Table 2). Except orkut, all these datasets come with input features that enable evaluating the accuracy of GNN training. The orkut dataset contains only the graph topology, so we only use it to measure epoch time for the scalability experiments. To run GNN training on it, we add synthetic input features of size 128, a train/val/test split of 60%/20%/20%. One out of two target classes is randomly assigned to each node.
@@
Representative GNN training systems. To compare across full-graph and mini-batch distributed GNN systems, we select representative systems that incorporate the main optimizations discussed in Section 2.2. We chose systems that have a stable open-source implementation, including implementations of our target models (GCN, GraphSage, and GAT) in their repositories [^2] DGL 1.1.2: https://github.com/dmlc/dgl; Quiver 0.1.0: https://github.com/quiver-team/torch-quiver; DistDGL: https://github.com/dmlc/dgl/tree/master/examples/distributed; PipeGCN: https://github.com/GATECH-EIC/PipeGCN; BNS-GCN: https://github.com/GATECH-EIC/BNS-GCN. We run all systems on top of PyTorch 2.0.1 and Python 3.8.10.
For full-graph training, we select PipeGCN (Wan et al., 2022b), which supports asynchronous training to overlap communication and computation, and BNS-GCN (Wan et al., 2022a), which uses sampling to reduce communication. Both systems partition the entire graph topology and features, and store each partition in the memory of one GPU.
For mini-batch training, we choose popular data-parallel systems that support GPU-based sampling, efficient input feature loading, and feature caching. Specifically, we use DGL (Wang et al., 2019), which only supports single-GPU caching, and Quiver (Tan et al., 2023), which supports distributed caching. DistDGL (Zheng et al., 2020) is built on top of DGL and performs mini-batch training on a distributed cluster of hosts. It only supports CPU-based training and does not support caching. All the mini-batch systems we consider support only synchronous training.
Model selection and implementations. We consider standard GNN models that are commonly used to evaluate training systems: GCN, GraphSage, and GAT. We use the implementations provided with the distributions of the systems we benchmark and validate that the semantics of the implementations are consistent across different systems. This is to make sure that the model implementations are fully compliant and optimized for the respective systems and avoid introducing biases. DGL and Quiver include implementations of all the three models in its distribution. The PipeGCN distribution includes GraphSAGE, while the GCN and GAT implementations for both PipeGCN and BNS-GCN are included in the BNS-GCN distribution. For mini-batch training, we considered the standard Neighborhood Sampling algorithm for our evaluation (Hamilton et al., 2017).
Hyperparameter search space. The search space for hyperparameters is shown in Table 3. The first set of hyperparameters define the architecture of the GNN models. Fanout is related to the neighborhood sampling algorithm used in mini-batch training. Finally, we consider general model training parameters such as learning rate, dropout and batch size. To search this parameter space, we use random search for all the parameters to obtain the best set of hyperparameters giving the highest validation accuracy.
@@ 4. Performance Evaluation
We evaluate the performance of our representative GNN training systems and answer two questions: Q1) How long do representative systems take to reach the same target accuracy? Q2) How do these systems scale with a varying number of GPUs and hosts? We mainly use time-to-accuracy as the main metric for performance comparison since it is a more informative metric than epoch time, as we will discuss.
We conduct end-to-end performance measurements because mini-batch and full-graph training systems have different implementations and use different optimizations. For a more in-depth discussion of the effect of specific systems optimizations, it is necessary to restrict the evaluation to one class of systems, as done in (Yuan et al., 2024) for mini-batch systems.
@@ 4.1. Time-to-Accuracy
We start by answering question Q1: How long do representative systems take to reach the same target accuracy?
@@
Methodology Time-to-accuracy is the time taken by any system to train to a target accuracy. To measure it, we select the target accuracies by leveraging hyperparameter tuning. Given a dataset and a GNN model, we first find a model architecture on which both training approaches (full-graph and mini-batch) converge to a similar and close-to-best accuracy, which we call convergence accuracy. We select the target accuracy as the minimum convergence accuracy across the two training approaches. We then use the same model architecture across all GNN training systems and measure their time to reach the target accuracy.
These experiments consider the two host types described in Section 3: a higher-end NVLink host and a lower-end PCIe host.
4.1.1. Single NVLink host
Table 4. Time to accuracy for GraphSAGE on a single NVLink host with 4 GPUs. Architectures (# layers, hidden size): pubmed - (3,256), ogbn-arxiv - (2,512), reddit - (3,1024), ogbn-products - (3,256), ogbn-papers100m - (2, 128). ET - Epoch time, TTA - Time to accuracy. For ogbn-papers100m DGL uses UVA mode for sampling and Quiver performs distributed caching. (Best results shown in bold.)
System pubmed ogbn-arxiv reddit ogbn-products ogbn-papers100m
Epoch time (s) PipeGCN (FG, Sync) 0.009 0.057 1.845 2.086 1.094
PipeGCN (FG, Async) 0.008 0.011 0.370 0.341 0.644
BNS-GCN (FG, Sync) 0.012 0.017 0.483 0.473 0.846
DGL (MB, Sync) 0.007 0.110 2.262 1.335 3.428
Quiver (MB, Sync) 0.005 0.131 1.994 2.117 1.878
Time to accuracy (s) PipeGCN (FG, Sync) 0.87 11.03 202.95 479.81 984.33
PipeGCN (FG, Async) 1.02 3.44 103.47 122.64 553.75
BNS-GCN (FG, Sync) 1.59 4.51 144.98 217.66 262.11
DGL (MB, Sync) 0.44 1.10 6.78 46.73 75.42
Quiver (MB, Sync) 0.36 0.65 9.97 84.70 41.31
Table 5. Time to accuracy with 3 NVLink hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached). (Best results shown in bold.)
System Time (s)
Epoch time (s) PipeGCN (FG, Sync) 1.139
PipeGCN (FG, Async) 0.352
BNS-GCN (FG, Sync) 0.432
DistDGL (MB, Sync) 4.452
Time to accuracy (s) PipeGCN (FG, Sync) 1093.63
PipeGCN (FG, Async) 306.59
BNS-GCN (FG, Sync) 133.92
DistDGL (MB, Sync) 35.62
Table 6. Time to accuracy and epoch time for GraphSAGE with a single PCIe host. Notation: ET - Epoch time, TTA - Time to accuracy. Architectures (# layers, hidden size): pubmed (3,256), ogbn-arxiv (2,512), reddit (4,1024) & ogbn-products (5,256). (Best results shown in bold.)
@@ 4.1.1. Single NVLink host
ET (s) PipeGCN (FG, Sync) 0.024 0.188 5.602 5.835
PipeGCN (FG, Async) 0.020 0.067 3.278 2.526
BNS-GCN (FG, Sync) 0.031 0.072 3.399 2.915
DGL (MB, Sync) 0.019 0.314 9.416 7.081
Quiver (MB, Sync) 0.019 0.216 10.272 9.246
TTA (s) PipeGCN (FG, Sync) 2.40 36.16 616.25 1341.95
PipeGCN (FG, Async) 2.35 20.75 917.80 909.43
BNS-GCN (FG, Sync) 2.91 19.33 1019.64 1340.80
DGL (MB, Sync) 1.25 3.14 28.25 247.83
Quiver (MB, Sync) 1.32 1.08 51.36 369.85
Table 7. Time to accuracy and epoch time with 3 PCIe hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached). (Best results shown in bold.)
System Time (s)
Epoch time (s) PipeGCN (FG, Sync) 22.00
PipeGCN (FG, Async) 1.03
BNS-GCN (FG, Sync) 2.46
DistDGL (MB, Sync) 38.37
Time to accuracy (s) PipeGCN (FG, Sync) 19141.48
PipeGCN (FG, Async) 956.34
BNS-GCN (FG, Sync) 859.26
DistDGL (MB, Sync) 767.38
Figure 2. Time to accuracy for ogbn-arxiv, reddit, ogbn-products for GraphSAGE (single NVLink host).
Figure 3. Time to accuracy for ogbn-papers100M on (3 NVLink hosts)
Table 4 shows the time to accuracy when training GraphSAGE on a single NVLink host. Full-graph training systems have a consistently lower epoch time (ET) than mini-batch training systems because they perform only one training iteration per epoch. However, their time-to-accuracy (TTA) is substantially higher compared to mini-batch systems. For example, on the largest graph we consider, ogbn-papers100m, Quiver is 6.4$\times$ faster than BNS-GCN. Similarly, for pubmed, Quiver is 2.4×2.4$\times$2.4 × faster than PipeGCN sync (lowest speed-up we observed) and for reddit, DGL is 15.2×15.2$\times$15.2 × faster than PipeGCN async (highest speed-up we observed). On average, the fastest mini-batch systems are 6.3×6.3$\times$6.3 × faster than the fastest full-graph system we considered for these datasets. This is because mini-batch training systems perform multiple training iterations and model updates per epoch and thus converge in fewer epochs.
Figure 2 shows the time-to-accuracy curves. The dotted line indicates the target accuracy. We observe mini-batch training consistently has a much faster time to accuracy across graphs with different sizes. Results on other models are shown in the extended version (Bajaj et al., 2024).
The experiments also show the impact of the systems optimizations introduced in previous work. Among the full-graph training systems, PipeGCN can substantially decrease epoch time and time-to-accuracy using asynchronous training compared to its baseline synchronous version. Sampling, as proposed by BNS-GCN, is less effective than asynchronous training on the smaller graphs, but it yields the fastest time-to-accuracy among all full-graph training systems on ogbn-papers100m.
All full-graph systems we consider fully cache the graph in GPU memory. Mini-batch training systems also cache all the graph in the GPU memory with the exception of ogbn-papers100m, which cannot be cached by DGL. This is because DGL only supports caching when the entire graph fits in the memory of each GPU, unlike Quiver which implements a distributed cache where the graph can be partitioned among multiple GPUs. Despite the lack of caching, DGL still achieves a lower time-to-accuracy than the full-graph training systems on ogbn-paper100m.
4.1.2. Multiple NVLink hosts
We perform a similar experiment using three NVLink hosts with the largest of the graphs we consider, ogbn-papers100m.
Table 5 shows the results. Figure 3 shows the corresponding time-to-accuracy curves. Quiver does not support multi-node training, so we do not include it in these experiments. We also replace DGL with its distributed version, DistDGL. The main trends are similar to the ones observed for the single-host experiments. DistDGL, which uses mini-batch training, has a much larger epoch time than the other systems, which use full-batch training. Its time-to-accuracy, however, is much shorter than the other systems. It is worth noting that DistDGL has a much larger time-to-accuracy than DGL on a single host (see Table 4). This is because DistDGL uses CPU-based sampling and does not support caching, unlike DGL.
4.1.3. PCIe hosts
@@
Table 6 reports the results for single-host training with GraphSAGE using a lower-end host setup, where NVLink is not available and cross-GPU communication within one host must occur over a slower PCIe bus. Compared to the NVLink experiments of Table 4, the epoch time of all systems is larger because the GPUs are less powerful. Full-graph training systems experience a stronger degradation in performance than DGL because they must run cross-GPU communication over PCIe, whereas DGL avoids communication by using data parallelism. Quiver also observes a performance degradation since loading data from other GPU’s cache is not as effective over the PCIe bus as it is over NVLink.
Table 7 futher compares all systems using 3 PCIe hosts for the ogbn-papers100m graph. In the PCIe setup, full-graph training systems need a multi-node setup to fully cache the ogbn-papers100m graph, which is larger than the others. In the setup with 3 PCIe hosts, hosts are connected by a slower network (1 GBps instead of 25 GBps). Like in the previous experiments, DistDGL has a lower time-to-accuracy than the other full-graph training systems, but the difference is much smaller than with the NVLink setup. This is because DistDGL requires cross-host communication to perform sampling and feature loading. These costs become bottlenecks over the slow 1GBps network of the PCIe host. By contrast, full-graph training systems are optimized to minimize communication and amortize its cost.
Takeaway. Mini-batch training enables faster time-to-accuracy compared to full-graph training since it requires fewer epochs to reach a target accuracy. This is because full-graph training systems update the model only once per epoch, whereas mini-batch training systems perform multiple iterations per epoch and update the model at each iteration. By updating the model more frequently, mini-batch training systems typically converge faster than full-graph training ones, even if their epoch time is longer.
These findings stress that time-to-accuracy measurements should always be included to provide a comprehensive comparison of systems using different training approaches. Comparing only in terms of epoch time, as is commonly done in prior work (see Table 1) does not provide a comprehensive comparison.

### v3

@@ 2.3. Motivation
The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field. A significant amount of work on GNN training do not compare or evaluate their systems against systems using a different training approach and training pipeline (full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021). Some prior work focuses only on the data-parallel mini-batching pipeline of Figure 1(b) and compares the performance of different low-level data management optimizations on top of the same underlying system (Yuan et al., 2024). In the following, we review 33 papers on GNN training that make claims comparing systems in the two classes. We summarize the claims in these papers in Table 1.
@@
We classify the papers into three broad groups. The first group consists of papers that propose full-graph training. Out of the 17 papers in this group, 10 provide experimental evidence. The rest 7 papers make claims that full-graph training is superior in terms of higher accuracy, lower epoch time, and convergence guarantees and support the claims by only citing other papers. The second group includes 12 papers that propose mini-batch training. Among these papers, only 4 provide experimental evidence to support claims that full-graph training is difficult to scale and requires more memory, while mini-batch training has lower time-to-accuracy. The last group consists of 4 papers that do not propose one specific training approach. Overall, we find the following shortcomings in the existing literature, which motivate our work.
@@
Conflicting claims about how the two approaches compare. Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems. Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level. Other works claim full-graph training is impractical, requires high memory, and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021). Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of the scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
@@
Comparing performance only in terms of epoch time. Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023). As we will show, a comparison in terms of time-to-accuracy is more informative given the differences in the two training approaches. Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
@@
Experiment design. Some works (Jia et al., 2020; Liu et al., 2021; Li et al., 2024; Kaler et al., 2022) do not always use the same GNN model (e.g., GraphSAGE or GAT) for the two approaches as a basis of comparison. This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique. Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023b, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022). As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa. Using hyperparameter tuning is necessary to perform a fair comparison.
@@
Table 2. Dataset statistics. (*) The input features and the training data for Orkut are synthetic.
Datasets Nodes Edges Classes Features Train/Val/Test
Pubmed 19k 108k 3 500 18k/0.5k/1k
Arxiv 169k 1.1M 40 128 91k/29k/48k
Reddit 232k 11.6M 41 602 152k/23k/55k
Products 2.4M 61.9M 47 100 196K/49K/2.2M
Orkut 3M 117M 2* 128* 1.8M/614k/614k*
Papers100M  [^1] M 1.6B 172 128 1.2M/125k/214k
3. Experimental methodology
@@
Datasets and models. Our experiments use six datasets: pubmed (Namata et al., 2012), ogbn-arxiv (Hu et al., 2020), reddit (Hamilton et al., 2017), ogbn-products (Hu et al., 2020), orkut (Yang and Leskovec, 2012) and ogbn-papers100m (Hu et al., 2020) (see Table 2). These datasets have varying average degrees, ranging from 5.6 to 60, and input feature sizes. They also differ on the fraction of vertices in the training set: in pubmed, this includes most vertices, whereas in papers it includes only 1.1% of the vertices. Except for orkut, all these datasets come with input features that enable evaluating the accuracy of GNN training. The orkut dataset contains only the graph topology, so we only use it to measure epoch time for the scalability experiments. We consider the three standard GNN models used in previous work: GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2017).
@@
Representative GNN training systems. To compare across full-graph and mini-batch distributed GNN systems, we select representative systems that have publicly available and stable implementations and incorporate the main design choices discussed in Section 2.2, as summarized in Figure 2. We run all systems on top of PyTorch 2.0.1 and Python 3.8.10.
For full-graph training, in the training data management dimension of Figure 2, our evaluation only considers the more favorable situation where the entire training data is fully cached in the memory of the GPUs since they can entirely avoid the overhead of GPU-host communication. All systems adopt a pipeline similar to the one of Figure 1(a). The Full-Graph baseline is synchronous and does not use any optimization that can impact accuracy (Wan et al., 2023a). In the dimension of algorithmic optimization, we consider three baselines representing three classes of algorithmic optimizations: PipeGCN for asynchronous training (Wan et al., 2022b, 2023a), BNS-GCN for sampling (Wan et al., 2022a, [n.d.]a), and AdaQP for message quantization (Wan et al., 2023c, [n.d.]b).
For mini-batch training, we consider three baselines: DGL (Wang et al., 2019; dgl, 2023), its distributed version DistDGL (Zheng et al., 2020; dgl, 2023), and Quiver (Tan et al., 2023; Quiver, 2023). These cover the system-level optimizations discussed in Section 2.2 (see Figure 2). All systems adopt a pipeline similar to the one of Figure 1(b). In terms of sampling implementations, DistDGL uses CPU-based sampling, while both DGL and Quiver support GPU-based sampling. DistDGL partitions the training dataset across multiple hosts. In terms of GPU caching, DGL only supports local GPU caching while Quiver also supports distributed GPU caching. Quiver uses distributed input feature caching for the Orkut and Papers100M datasets, which are too large for local caching. Mini-batch training can use different sampling algorithms to select a subset of the k-hop neighbors of the vertices in the mini-batch. We consider the standard Neighborhood Sampling (NS) algorithm, which is available for both DGL and Quiver, ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
Hyperparameter search. We run an extensive hyperparameter tuning. We validated our search by matching the best test accuracies we found in the literature on GNN training systems we reviewed. To manage the large search space, we iterate multiple times over three phases, where we tune a set of hyperparameters and fix all others until we don’t observe any more improvements in accuracy. In the first phase, we use grid search for the architectural hyperparameters, namely the number of layers or the hidden feature size, which have a limited number of discrete values. In the second phase, we run a random search for the remaining hyperparameters, such as the learning rate. Mini-batch training requires tuning the micro-batch size and additional sampling hyperparameters compared to full-graph training. We tune these in a third phase using Bayesian optimization, reducing the search space by considering only a few commonly used discrete values. In the appendix, we report our hyperparameter space in table 11 and show that mini-batch training is less sensitive than full-graph training to the tuning of the architectural hyperparameters.
@@ 4. Performance Evaluation
We evaluate the performance of our representative GNN training systems and answer two questions: Q1) How long do representative systems take to reach the same target accuracy? (Section 4.1), and Q2) How do these systems scale with a varying number of GPUs and hosts? (Section 4.2).
@@ 4.1. Time-to-Accuracy
Methodology Time-to-accuracy is the time taken by any system to train to a target accuracy. To measure it, we select the target accuracies by leveraging hyperparameter tuning. Given a dataset and a GNN model, we first find a model architecture on which both training approaches (full-graph and mini-batch) converge to a similar and close-to-best accuracy, which we call convergence accuracy. We select the target accuracy as the minimum convergence accuracy across the two training approaches. We then use the same model architecture across all GNN training systems and measure their time to reach the target accuracy. The time-to-accuracy considers only the final training phase and not the hyperparameter search. For mini-batch systems, we consider the common Neighborhood Sampling algorithm as default because it is available on all baselines and it consistently achieves an accuracy close to the full-graph training methods. We report the hyperparameters in the appendix Appendix. These experiments consider the two host types described in Section 3: a higher-end NVLink host and a lower-end PCIe host.
Table 3. Time to accuracy (TTA) and epoch time (ET) for GraphSAGE (NVLink host).
System pubmed ogbn- reddit ogbn- ogbn-
arxiv products papers100m
ET (s) Full-Graph 0.009 0.057 1.845 2.086 1.094
PipeGCN 0.008 0.011 0.370 0.341 0.644
BNS-GCN 0.012 0.017 0.483 0.473 0.846
AdaQP 0.027 0.029 0.830 0.312 0.471
DGL 0.007 0.110 2.262 1.335 3.428
Quiver 0.005 0.131 1.994 2.117 1.878
TTA (s) Full-Graph 0.87 11.03 202.95 479.81 984.33
PipeGCN 1.02 3.44 103.47 122.64 553.75
BNS-GCN 1.59 4.51 144.98 217.66 262.11
AdaQP 0.81 19.86 41.5 53.04 160.41
DGL 0.44 0.55 6.78 40.05 75.42
Quiver 0.33 0.65 5.98 63.52 41.31
Figure 3. Convergence curve for GraphSAGE (NVLink host).
NVLink host. Table 3 and Figure 3 show the time to accuracy when training GraphSAGE on a single NVLink host with 4 GPUs. Full-graph training systems have a consistently lower epoch time (ET) than mini-batch training systems because they perform only one training iteration per epoch. However, their time-to-accuracy (TTA) is substantially higher compared to mini-batch systems. For example, on the largest graph we consider, ogbn-papers100m, Quiver is 3.9×3.9$\times$3.9 × faster than AdaQP. Similarly, for pubmed, Quiver is 2.5×2.5$\times$2.5 × faster than AdaQP; for arxiv, DGL is 6.3×6.3$\times$6.3 × faster than PipeGCN; for reddit, Quiver is 6.9×6.9$\times$6.9 × faster than AdaQP (highest speed-up we observed); for products, DGL is 33% faster than AdaQP (lowest speed-up we observed). Results on other models are shown in the appendix in table 17.
Comparing the performance of systems in the same class shows the impact of the systems optimizations introduced in previous work (see Figure 2). All full-graph systems we consider benefit from fully caching the graph in GPU memory. PipeGCN can substantially decrease ET and TTA compared to Full-Graph. By using asynchronous training and overlapping communication with computation, PipeGCN can mitigate the cost of synchronous cross-GPU vertex feature communication at each layer, which is high even when NVLink. Boundary-node sampling, as proposed by BNS-GCN, is less effective than asynchronous training on the smaller graphs, but it yields the fastest TTA among all full-graph training systems on ogbn-papers100M. The use of of synchronous communication in BNS-GCN results in larger ET than PipeGCN for all datasets, but using fresh training data contributes to achieving a shorter number of epochs to convergence with papers100M. AdaQP also combines synchrony with reduced communication costs thanks to message quantization. It speeds up ET significantly compared to Full-Graph, but not compared to other full-graph optimizations. Nonetheless, AdaQP can have faster TTA than other full-graph graph systems, since for some datasets it requires fewer epochs to converge.
Among mini-batch systems, DGL supports local full caching and can cache all the graphs in Table 3 except papers100m. When full local caching is possible, DGL does not need to transfer input vertex features to the GPUs, maximizing GPU utilization. Quiver also uses full local caching, achieving a similar ET and TTA. However, Quiver implements a partial distributed cache across multiple GPUs that can also cache ogbn-papers100m. With distributed caching, some input vertex features may still be transferred among GPUs using NVLink buses, which is cheaper than loading the features from the host memory over the slower PCIe bus. This is why Quiver has a much lower time-to-accuracy than DGL for papers100M.
Table 4. Time to accuracy (TTA) and epoch time (ET) for GraphSAGE (PCIe host).
@@
ET (s) Full-Graph 0.024 0.188 5.602 5.835
PipeGCN 0.020 0.067 3.278 2.526
BNS-GCN 0.031 0.072 3.399 2.915
DGL 0.019 0.314 9.416 7.081
Quiver 0.019 0.216 10.272 9.246
TTA (s) Full-Graph 2.40 36.16 616.25 1341.95
PipeGCN 2.35 20.75 917.80 909.43
BNS-GCN 2.91 19.33 1019.64 1340.80
DGL 1.25 1.57 28.25 247.83
Quiver 1.22 1.08 30.82 323.61
@@
PCIe host. Table 4 reports the results for single-host training with GraphSAGE using a lower-end host setup, where NVLink is not available and all cross-GPU communication must occur over a slower PCIe bus. The 4 GPUs are computationally less powerful and have less memory. The papers100M graph is too large to be fully cached in a single host by our full-graph training systems. We cannot run AdaQP since its implementation is not compatible with our lower-end GPUs.
Like in the previous hardware setups, mini-batch training systems show a larger ET (with the exception of pubmed) but a shorter TTA on all datasets. Full-graph training systems must run cross-GPU vertex feature exchange over the slow PCIe bus. Optimizations that reduce the communication cost, like asynchrony (PipeGCN) and sampling (BNS-GCN), can reduce TTA, but the cost remains high. DGL and Quiver avoid vertex feature communication by replicating the training data on all GPUs so the advantage over the full-graph training systems is larger than with the NVLink host, reaching 32×32$\times$32 × for Reddit.
The impact of dataset characteristics. Different datasets have an impact on the performance of different systems. While the impact on TTA is hard to understand due to the stochastic nature of training, the impact on ET can be more easily analyzed.
The communication cost of full-graph training systems is determined by the number of boundary vertices across partitions. Reddit and products, for example, have a large number of boundary vertices, so the ET of the Full-Graph baseline is much larger compared to DGL and Quiver. Optimizations that reduce this communication cost used by PipeGCN and BNS-GCN can reduce this gap significantly. For mini-batch training, the training cost is determined by the size of the micro-batches assigned to each GPU. Datasets that require a larger number of GNN layers, such as products and reddit which requires 5 and 4 layers respectively, result in a larger ET for mini-batch training systems compared to the other full-graph training systems. In contrast, papers100M only requires 2 layers so the gap in ET is smaller.
Different mini-batch sampling algorithms. Finally, we combine DGL with different mini-batch sampling algorithms than Neighborhood Sampling (NS).ClusterGCN achieves on average 2.8×2.8$\times$2.8 × faster TTA than NS across datasets, ranging from 0.14×0.14$\times$0.14 × slower to 12.3×12.3$\times$12.3 × faster depending on the dataset, and its ET is 6.4×6.4$\times$6.4 × faster on average. GraphSaint has 3.2×3.2$\times$3.2 × faster TTA than NS on average, ranging from 0.5×0.5$\times$0.5 × slower to 16.1×16.1$\times$16.1 ×, and faster ET by 9.3×9.3$\times$9.3 × on average (see appendix tables 24 and 25).
Takeaway. Mini-batch training systems converge faster than full-graph training ones because they require fewer epochs to reach a target accuracy, despite having longer epochs. We observe this across all models, datasets, and hardware configurations we consider. Avoiding data transfers by caching input data in GPU memory has a strong impact on the performance of mini-batch training. For full-graph training, asynchrony (PipeGCN) and sampling (BNS-GCN) can reduce communication cost and thus ET, resulting in faster TTA. AdaQP can achieve even faster TTA even if its ET is larger than other full-graph methods, showing the importance of using algorithmic optimizations besides system-level optimizations. For mini-batch training, using different sampling algorithms can speed up TTA significantly in many cases.